OpenStudy (anonymous):

b + 6.7 = 19.4 Will give medal!!!!!!!!!!

OpenStudy (anonymous):

b+6.7=19.4 we need to get b all alone, so we subtract 6.7 from both sides b=12.7

OpenStudy (anonymous):

i agree also u can check ur work by plugging the b value u get into the equation 12.7 + 6.7 = 19.4

OpenStudy (anonymous):

oh thx

OpenStudy (anonymous):

wait can you guys help me whith a few more?

OpenStudy (anonymous):

with

OpenStudy (anonymous):

sure no problem

OpenStudy (anonymous):

oh thx

OpenStudy (anonymous):

Solve for f. f + (–15) = –32 A. 47 B. 17 C. –47 D. –17

OpenStudy (anonymous):

f = 15 - 32 so f = -17

OpenStudy (anonymous):

Marcus biked 15.7 miles on Saturday and 13.9 miles the following Thursday. Ellen biked the same total distance as Marcus, but only biked 12.8 miles on Saturday. How far did Ellen bike on Thursday? A. 16.8 miles B. 29.6 miles C. 2.9 miles D. 28.5 miles

OpenStudy (anonymous):

12.8+x=15.7+13.9 12.8+x=29.6 x=16.8

OpenStudy (anonymous):

Which phrase correctly describes the location of –52 on the number line? A. to the left of –67 B. to the left of –38 C. to the right of –15 D. to the right of 50

OpenStudy (anonymous):

yup agreed on x = 16.8

OpenStudy (anonymous):

B. to the left of -38

OpenStudy (anonymous):

What is the perimeter of an isosceles triangle with base of 13 m and congruent sides of 9 m?

OpenStudy (anonymous):

i think it's 32 .... 9 + 9 + 13 = 32

OpenStudy (raffle_snaffle):

B

OpenStudy (anonymous):

Anne walked 3.2 kilometers. How many meters did Anne walk? A. 3200 meters B. 0.0032 meters C. 320 meters D. 0.032 meters

OpenStudy (raffle_snaffle):

Now I want my medal

OpenStudy (anonymous):

i think since it said meters should be 3.2 *100 = 320

OpenStudy (anonymous):

The perimeter of this rectangle can be calculated using the equation P = 2 × 18.5 + 2 × 7.2. Which other equation could be used to calculate the perimeter? A. P = 2 (18.5 + 7.2) B. P = 2 × 18.5 + 7.2 C. P = (2 • 18.5) + 7.2 D. P = 2 × 18.5 × 7.2

OpenStudy (anonymous):

gf

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

bk

OpenStudy (anonymous):

i think i'm wrong on the above meters question so i'll look into it again

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

A works for the above problem 2(18.5 +7.2) is equal to 2*18.5 + 2*7.2... they both equal 51.4 by checking that they are the same

OpenStudy (anonymous):

last one fiber

OpenStudy (anonymous):

Which phrase correctly orders –34, 15, –23, and 2? A. –34 < –23 < 2 < 15 B. 2 < 15 < –34 < –23 C. 2 < 15 < –23 < –34 D. –34 < –23 < 15 < 2

OpenStudy (anonymous):

i think kilometers 3.2 *1000 = 3200

OpenStudy (anonymous):

Are you sure fiber

OpenStudy (anonymous):

should be A - 34 ,-23, 2, 15

OpenStudy (anonymous):

Fib

OpenStudy (anonymous):

I am in love with you lol jk but 86% i passed

OpenStudy (anonymous):

that how i would do it but i'm not certain on the kilometer question

OpenStudy (anonymous):

sorry if i cost u 14 percent

OpenStudy (anonymous):

No as long as i passed i am happy and your my new best friend haha

OpenStudy (anonymous):

lol thxs glad u passed hope ur able to understand it as well

OpenStudy (anonymous):

I do wait are you a dude or a girl?

OpenStudy (anonymous):

i'm male

OpenStudy (anonymous):

Yay

OpenStudy (anonymous):

Dude bros lol

OpenStudy (anonymous):

Guy bros

OpenStudy (anonymous):

yeah well i gotta go grad some food laterz hollar back if u need more help

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well thx a ton guys

OpenStudy (anonymous):

i gtg guys bye all

OpenStudy (anonymous):

thx again fiber